OpenStudy (anonymous):

what is the inverse function of f(x)=x2-4 1) x2+4 2) sqrt(x+4) 3) sqrt(x)+4 4) sqrt(x-4)

OpenStudy (anonymous):

f(x)=x^2-4, y=x^2-4, x=y^2-4, x+4=y^2, sqrt(x+4)=y so the answer is f(x)=sqrt(x+4) HI AGAIn

OpenStudy (anonymous):

@Nnesha can you check him

OpenStudy (anonymous):

It is \[\sqrt{x+4}\]

OpenStudy (anonymous):

yes so we are correct

OpenStudy (freckles):

I hope the teacher gave some restriction to domain because the inverse isn't actually a function it is relation since you have y^2=x+4 which implies y=sqrt(x+4) or -sqrt(x+4)

OpenStudy (anonymous):

the onkly abswer can be the second

OpenStudy (anonymous):

I think you are right, but my teacher never says restrivtions

OpenStudy (freckles):

that's a bit odd

OpenStudy (anonymous):

yes it is

OpenStudy (anonymous):

@freckles In my online algebra 2 they never mentioned restrictions, but in precalc they do.

OpenStudy (anonymous):

@EmmaMink

OpenStudy (anonymous):

FLVS criminal idiot s wrote this i bet $20

OpenStudy (freckles):

inverse function of f(x)=x^2-4 doesn't exist unless they restricted the domain ex.f(x)=x^2-4 on [0,inf) has inverse g(x)=sqrt(x+4) ex. f(x)=x^2-4 on (-inf,0] has inverse g(x)=-sqrt(x+4) if they wanted an inverse relation you could say + or - sqrt(x+4)

OpenStudy (anonymous):

I'm a criinal idiot?

OpenStudy (anonymous):

they don't know any math, just play a trick on unsuspecting florida students

OpenStudy (anonymous):

forget I'm from russia

OpenStudy (anonymous):

@ttb123456789 no i am speaking of the FLVS

OpenStudy (freckles):

satellite just really hates the school that makes these questions

OpenStudy (freckles):

because it is not mathematically correct

OpenStudy (anonymous):

course i could be wrong

OpenStudy (anonymous):

i keep asking for question ID numbers from the wrong questions written by FLVS

OpenStudy (anonymous):

so i can bring them to their attention no one seems to want to provide them

OpenStudy (anonymous):

i'm confused

OpenStudy (emmamink):

dang ttb123456789 for only 11, you're more advanced than I am in math!